Can someone show me the y intercept and slope of these questons ?
1)6x+3y=54
2)4x+2y

Answer by Earlsdon(6100)

Yes on number 1
No on number 2, why, because you don't have an equation, you have an expression. To show the slope and y-intercept, you must have an equation, which means 4x+2y must be equal to something.
1) Rewrite the equation in the slope-intercept form: y = mx + b
6x + 3y = 54 Subtract 6x from both sides of the equation.
3y = -6x + 54 Divide both sides by 3.
y = -2x + 18, compare this with:
y = mx + b
m = -2 This is the slope.
b = 18 This is the y-intercept.